1/5 dilution of sunstance requires an additional 1/2 dilution final measurment of substances in diultion is 5.0 mg/mL. What is the concentration of the undiluted sample?
Answer #1

First dilution = 1/5

Second dilution = 1/2

So, total dilution

= first diultion x second dilution

= 1/5 x 1/2

= 1/10

So, dilution factor = 10

Final concentration = 5.0mg/mL

Concentration of the undiluted sample

= final concentration x dilution factor

= 5.0 x 10

= 50 mg/mL

Comment if any problem